# PriorRG: Prior-Guided Contrastive Pre-training and Coarse-to-Fine Decoding for Chest X-ray Report Generation

### v1

@@ 4.1 Experimental Settings
Datasets. (1) MIMIC-CXR (Johnson et al. 2019) is a large-scale, publicly available dataset of chest X-rays paired with free-text radiology reports. For each case, we organize data by “study id” and retrieve the most recent prior image when available. (2) MIMIC-ABN (Ni et al. 2020) is a curated subset of MIMIC-CXR that focuses exclusively on abnormal findings in reports. Following previous studies (Chen et al. 2020; Liu et al. 2024b; Xiao et al. 2025), we treat only the “Findings” section as the radiologist-generated report, discarding empty or non-informative entries. All experiments adhere to the official dataset partitions. Dataset statistics are presented in Table 6.
Evaluation metrics. Following protocols in (Wang et al. 2023a; Liang et al. 2024; Liu et al. 2024a), we evaluate the quality of AI-generated reports using both natural language generation (NLG) and clinical efficacy (CE) metrics. For NLG metrics, which measure linguistic similarities between AI- and radiologist-generated reports, we use BLEU-n (B-n), METEOR (MTR), and ROUGE-L (R-L). For CE metrics, we assess clinical relevance by computing the micro-average Precision (P), Recall (R), and F1-score (F1) across 14 observations annotated by CheXpert (Irvin et al. 2019).
Implementation details. The unified feature dimension is set to $d=768$, the number of latents to $N=128$, the maximum generation length to $K=100$ tokens, and the beam size to 3 during inference. We apply early stopping and learning rate scheduling for training stability. Additional implementation details are presented in Appendix A.
@@ 4.2 Main Results
Baseline approaches. We compare our PriorRG against 12 state-of-the-art (SOTA) methods, categorized into eight groups: knowledge graph-based methods (KiUT (Huang, Zhang, and Zhang 2023) and METransformer (Wang et al. 2023a)), contrastive learning framework (CoFE (Li et al. 2025b)), retrieval-augmented method (DCG (Liang et al. 2024)), memory alignment approaches (CMN (Chen et al. 2021) and MAN (Shen et al. 2024)), LLM-based approaches (R2GenGPT (Wang et al. 2023b), Med-LLM (Liu et al. 2024d), and R2-LLM (Liu et al. 2024a)), clinical context-driven model (SEI (Liu et al. 2024b)), longitudinal data-based method (HERGen (Wang, Du, and Yu 2025)), and human preference optimization (MPO (Xiao et al. 2025)).
Observation % SEI (Liu et al. 2024b) PriorRG (Ours)
@@
Comparison with SOTA methods. Table 1 presents a comparison between our PriorRG and SOTA methods on MIMIC-CXR (Johnson et al. 2019) and MIMIC-ABN (Ni et al. 2020) datasets. The results show that our PriorRG consistently surpasses existing methods in both linguistic quality and clinical accuracy. On the MIMIC-CXR dataset, PriorRG achieves a notable 3.6% improvement in B-4, indicating better matching accuracy in longer n-gram sequences. Additionally, gains in MTR (+1.3%) and R-L (+1.5%) highlight improvements in lexical diversity, synonym matching, and semantic consistency. In terms of CE metrics, PriorRG boosts the F1 by 3.8%, demonstrating greater reliability in clinical accuracy. On the MIMIC-ABN dataset, PriorRG achieves a 5.9% increase in B-1 and a 1.1% improvement in F1, emphasizing its effectiveness in describing abnormal findings. These improvements stem from our proposed prior-aware coarse-to-fine decoding, which enables the model to recognize individual clinical contexts and capture disease progression.
To comprehensively evaluate the generalization ability of our PriorRG in cross-dataset settings, we conduct experiments on the IU X-ray (Demner-Fushman et al. 2016) dataset under both zero-shot and supervised fine-tuning scenarios. The corresponding results are reported in Table 7. Our method consistently outperforms baselines across both settings, demonstrating strong generalization to unseen data distributions. In addition, we compare our PriorRG model with a recent representative baseline, MLRG (Liu et al. 2025a), along three key dimensions: report generation quality, inference time, and GPU memory usage. As shown in Table 10, PriorRG achieves superior performance across all evaluation metrics—indicating enhanced clinical and linguistic quality—while also exhibiting greater computational efficiency, with reduced inference time and memory consumption. These results further highlight the robustness and practicality of our method for real-world deployment.
Clinical accuracy of 14 observations. Table 2 presents the clinical accuracy of 14 observations extracted by the CheXpert (Irvin et al. 2019; Smit et al. 2020). Compared to SEI (Liu et al. 2024b), PriorRG achieves higher recall (R) in 12 out of 14 observations, with particularly significant improvements in Pneumonia, Pneumothorax, and Fractures—rare but clinically important conditions. This demonstrates that PriorRG reduces the risk of missed diagnoses and enhances its ability to detect low-frequency yet critical diseases, even without explicitly addressing class imbalance. Additionally, PriorRG outperforms SEI in F1-score (F1) for 13 out of 14 observations, highlighting its ability to reduce false positives and minimize false negatives, thereby improving the overall clinical reliability of generated reports.
Model Stage 1 Stage 2 NLG Metrics $\uparrow$ CE Metrics $\uparrow$
$C$ PI Hidden States B-1 B-2 B-3 B-4 MTR R-L P R F1
(a) 0.355 0.219 0.149 0.108 0.154 0.262 0.521 0.431 0.472
(b) 0.357 0.216 0.143 0.102 0.154 0.255 0.513 0.480 0.496
(c) 0.405 0.283 0.214 0.170 0.186 0.320 0.517 0.460 0.487
(d) 0.404 0.285 0.217 0.173 0.187 0.328 0.540 0.477 0.507
(e) 0.400 0.282 0.214 0.171 0.186 0.323 0.539 0.465 0.499
(f) 0.387 0.271 0.206 0.165 0.180 0.320 0.526 0.408 0.459
PriorRG 0.412 0.290 0.220 0.175 0.189 0.324 0.541 0.485 0.511
Evaluation with a large language model. We assess clinical consistency and expert alignment of generated reports using two metrics: “#Matched Findings” and the GREEN score (Ostmeier et al. 2024). The GREEN score combines clinical significance errors with “#Matched Findings,” offering a robust evaluation aligned with expert preferences. Both metrics are derived via the pre-trained GREEN-RadLlama2-7B model. We benchmark PriorRG against CMN (Chen et al. 2021), CvT2DistilGPT2 (Nicolson, Dowling, and Koopman 2023), SEI (Liu et al. 2024b), and MLRG (Liu et al. 2025a). Figure 4 shows that PriorRG significantly outperforms all baselines across both metrics, confirming its ability to generate clinically reliable, expert-aligned reports.
4.3 Ablation Study
Effect of prior-guided contrastive pre-training (Stage 1). We evaluate Stage 1 using a medical image-text retrieval task on the MIMIC-5x200 dataset (Johnson et al. 2019). Following the protocols in (Zhang et al. 2022; Zhou et al. 2024a), we construct the MIMIC-5x200 dataset by sampling 200 test set instances for each of five common diseases— Atelectasis, Cardiomegaly, Consolidation, Edema, and Pleural Effusion—from MIMIC-CXR, resulting in a total of 1,000 samples. Given a query image, we retrieve the top-K most similar reports from this dataset and evaluate performance using Category-Precision@K (Cat-P@K) and Study-Precision@K (Stu-P@K). Cat-P@K measures whether retrieved reports belong to the same disease category as the query image, while Stu-P@K evaluates whether retrieved reports originate from the same study. As illustrated in Figure 5, PriorRG surpasses BiomedCLIP (Zhang et al. 2023a), BioViL-T (Bannur et al. 2023), and ablated variants across both metrics. These results indicate that Stage 1 effectively encodes prior knowledge into semantically rich multimodal representations. Furthermore, Table 3 shows that removing Stage 1 (variant (f)) leads to significant drops in both clinical accuracy and language quality, confirming its central role in improving downstream report generation.
Metric w/ PI w/o PI w/ CC w/o CC w/ PK w/o PK
B-2 $\uparrow$ 0.289 0.288 0.294 0.139 0.294 0.140
F1 $\uparrow$ 0.516 0.514 0.508 0.492 0.512 0.497
Figure 6: Qualitative comparison of reports from the baseline (Liu et al. 2024b), our proposed PriorRG, and radiologists. Sentences in the radiologist-generated report are color-coded to match corresponding descriptions in the AI-generated reports. In PriorRG’s output, correct disease progression patterns and failure descriptions are marked in bold and underlined, respectively.
Model NLG Metrics $\uparrow$ CE Metrics $\uparrow$
B-2 B-4 MTR P R F1
LastOnly 0.278 0.168 0.185 0.554 0.466 0.506
Fine2coarse 0.284 0.172 0.187 0.549 0.482 0.514
PriorRG 0.290 0.175 0.189 0.541 0.485 0.511
Qualitative analysis. Figure 6 presents a comparative visualization of generated reports on the MIMIC-CXR dataset. In AI-generated reports, more diverse colors indicate broader coverage of clinical findings, while longer bars reflect more accurate and detailed descriptions. The results show that: (1) PriorRG generates high-quality drafts requiring minimal revision—for example, in Case 1, only a brief note on pleural effusion is added by the radiologist. (2) PriorRG captures disease progression effectively, as seen in the correct description of unchanged cardiomegaly in Case 1. (3) PriorRG responds to clinical context—e.g., in Case 2, the model correctly addresses concern about Dobbhoff tube placement and identifies potential complications. These examples demonstrate PriorRG’s ability to track disease progression and understand clinical intent. Additional examples and failure case analysis are provided in the Appendix.
Conclusion
We introduced PriorRG, a chest X-ray report generation framework that leverages clinical priors to enhance both image-text alignment and report quality. Specifically, we proposed a prior-guided contrastive pre-training scheme that mirrors diagnostic reasoning by using clinical context to guide spatiotemporal features extraction. Next, we presented a prior-aware coarse-to-fine decoding that progressively integrates clinical context, disease progression patterns, and hierarchical visual cues, enhancing the clinical accuracy and fluency of generated reports. Extensive experiments validated the effectiveness of our PriorRG on medical image-text retrieval and report generation. Future work will explore the organ-aware diagnosis framework (Gu et al. 2024) to further enhance interpretability.
Acknowledgments
The work was jointly supported by the National Natural Science Foundations of China [grant number: 62272364], the Provincial Key Research and Development Program of Shaanxi [grant number: 2024GH-ZDXM-47], the Research Project on Higher Education Teaching Reform of Shaanxi Province [grant number: 23JG003], the Fundamental Research Funds for the Central Universities and the Innovation Fund of Xidian University [grant number: YJSJ25012], and the High-Performance Computing Platform of Xidian University.
@@ Acknowledgments
Table 6: Statistics of the MIMIC-CXR (Johnson et al. 2019), MIMIC-ABN (Ni et al. 2020), and IU X-ray (Demner-Fushman et al. 2016) for training, validation, and test sets. “#Image”, “#Sample”, and “#Report” denote the number of medical images, samples, and radiology reports, respectively. “%PI”, “%Indication”, and “%History” represent the proportions of prior images, indications, and medical history, respectively.
[…]
Zero-shot Report Generation
R2GenGPT Meta-Radio’23 0.112 0.177
Med-LLM MM’24 0.125 0.185
PriorRG (Ours) - 0.178 0.211
Supervised Report Generation
METransformer CVPR’23 0.172 0.192
MAN AAAI’24 0.170 0.213
DCG MM’24 0.186 0.211
Med-LLM MM’24 0.168 0.209
R2-LLM AAAI’24 0.184 0.208
PriorRG (Ours) - 0.196 0.216
Table 7: Performance comparison of report generation on the IU X-ray dataset (Demner-Fushman et al. 2016) under zero-shot and supervised settings. Best results are highlighted in Bold.
Appendix A Implementation Details
Model architecture. The image encoder (RAD-DINO) (Pérez-García et al. 2024) remains frozen throughout training. In contrast, the text encoder (CXR-BERT) (Boecking et al. 2022), the report generator (DistilGPT2) (Sanh et al. 2020), projection heads, the attention-enhanced layer fusion network (ALF), and the spatiotemporal fusion network (STF) are fine-tuned during both training stages. All experiments are conducted on a single NVIDIA V100 GPU, with a memory usage of approximately 26 GB.
Training schedule. We adopt a two-stage training pipeline. For the MIMIC-CXR dataset (Johnson et al. 2019), the model is trained for 30 epochs in Stage 1 (prior-guided contrastive pre-training) with a batch size of 32, followed by another 30 epochs of fine-tuning in Stage 2 (prior-aware coarse-to-fine decoding) with a batch size of 16. In Stage 1, the learning rate is set to 5e-5. In Stage 2, the report generator is optimized with a learning rate of 5e-5, while all other components are trained with a learning rate of 5e-6. For the MIMIC-ABN dataset (Ni et al. 2020), as a derivative of MIMIC-CXR, the model is initialized with weights pre-trained on MIMIC-CXR and fine-tuned for 30 epochs using a learning rate of 5e-6 and a batch size of 16.
Metrics implementation details. All metrics are implemented using the pycocoevalcap (Chen et al. 2015) and f1chexbert (Smit et al. 2020) libraries. For all metrics, higher values indicate better performance.
Medical image-text retrieval. BioViL-T [^1] https://huggingface.co/microsoft/BiomedVLP-BioViL-T (Bannur et al. 2023) and BiomedCLIP [^2] https://huggingface.co/microsoft/BiomedCLIP-PubMedBERT˙256-vit˙base˙patch16˙224 (Zhang et al. 2023a) are implemented using their officially released pre-trained weights and inference pipelines. Our PriorRG model is initialized with weights obtained from Stage 1 pre-training. For all three methods, the medical image-report retrieval is conducted by calculating feature similarity between image and report embeddings, followed by retrieving the top-K most similar reports.
Appendix B Additional Experiments
Method M/S Stage 1 Stage 2
Trait #Param Trait #Param
MLRG M DP 191M DP+$C$ 296M
PriorRG S DP+$C$ 239M DP+$C$ 335M
Table 8: Comparison between our PriorRG and the recent baseline MLRG (Liu et al. 2025a) in terms of input data and parameter count. M/S: multi-view or single-view current images; “CC” and “DP” indicate whether a method models clinical context and disease progression, respectively.
Method Stu-P@1 Stu-P@3 Stu-P@5 Cat-P@1 Cat-P@3 Cat-P@5
MLRG 0.096 0.096 0.088 0.450 0.448 0.437
PriorRG w/o PK 0.279 0.238 0.191 0.601 0.574 0.549
PriorRG w/o CC 0.280 0.225 0.173 0.602 0.561 0.526
PriorRG w/o PI 0.393 0.324 0.245 0.637 0.597 0.548
PriorRG (Ours) 0.402 0.334 0.251 0.640 0.596 0.544
Method Venu NLG Metrics $\uparrow$ CE Metrics $\uparrow$
Time
(s/sample)
$\downarrow$
GPU $\downarrow$
B-1 B-2 B-3 B-4 MTR R-L P R F1 RG
Table 10: Performance comparison of report generation on the MIMIC-CXR dataset. Best results are highlighted in Bold. RG denotes F1-RadGraph (Jain et al. 2021) metric, which evaluates the overlap of clinical entities and relations, and better aligns with radiologists’ assessments (Yu et al. 2023).
B.1 Comparison with SOTA Methods on the IU X-ray Dataset in Zero-shot and Supervised Settings
Following previous works (Chen et al. 2020, 2021), we split the IU X-ray dataset (Demner-Fushman et al. 2016) into training, validation, and test sets with a 7:1:2 ratio. The detailed distribution is presented in Table 6.
Results in Table 7 show the following: (1) In the zero-shot setting, PriorRG achieves the highest BLEU-4 (0.178) and METEOR (0.211), clearly outperforming baselines like Med-LLM (Liu et al. 2024d) and R2GenGPT (Wang et al. 2023b). (2) In the supervised setting, PriorRG again achieves the best results, with BLEU-4 of 0.196 and METEOR of 0.216, outperforming strong baselines including METransformer (Wang et al. 2023a), MAN (Shen et al. 2024), DCG (Liang et al. 2024), Med-LLM (Liu et al. 2024d), and R2-LLM (Liu et al. 2024a). (3) These results demonstrate the strong generalization capability of our PriorRG across datasets and its robustness, even when prior images are unavailable.
B.2 Comparison with Concurrent Work MLRG on the MIMIC-CXR Dataset
MLRG (Liu et al. 2025a), a concurrent work accepted at CVPR 2025, also jointly models temporal visual changes and clinical context. To enable a comprehensive comparison, we evaluate both methods in terms of methodological design (Table 8), medical image-text retrieval (Table 9), and report generation (Table 10). (1) Methodology: Unlike MLRG (Liu et al. 2025a), our PriorRG explicitly models disease progression and clinical context during cross-modal alignment (Stage 1; see Table 8). This design leads to significantly improved retrieval performance, as reflected by higher Stu-P@K and Cat-P@K scores (Table 9). (2) Report Generation: On the chest X-ray report generation task, PriorRG consistently outperforms MLRG across natural language generation (NLG) and clinical efficacy (CE) metrics, achieving a +1.7% improvement in BLEU-4 (B-4) and +0.6% in F1-CheXbert (F1). (3) Efficiency and Deployment: PriorRG requires only a single-view current image while delivering higher-quality reports. It achieves over 2× faster inference and reduces GPU memory usage by nearly 5 GB. All comparisons are conducted on the MIMIC-CXR test set (3,852 samples) under identical hardware, batch size, and num_worker configurations, demonstrating PriorRG’s practical efficiency and deployment readiness.
B.3 Qualitative Analysis

### v2

@@ 4.1 Experimental Settings
Datasets. (1) MIMIC-CXR (johnson-mimic-cxr-jpg) is a large-scale, publicly available dataset of chest X-rays paired with free-text radiology reports. For each case, we organize data by “study id” and retrieve the most recent prior image when available. (2) MIMIC-ABN (mimic-abn-ori) is a curated subset of MIMIC-CXR that focuses exclusively on abnormal findings in reports. Following previous studies (chen-etal-2020-generating; sei; aaai-2025-mpo), we treat only the “Findings” section as the radiologist-generated report, discarding empty or non-informative entries. All experiments adhere to the official dataset partitions. Dataset statistics are presented in Appendix Table 11.
Evaluation metrics. Following protocols in (wang2023metransformer; liang2024divide-acmmm-24; aaai-liu2024bootstrapping-llm), we evaluate the quality of AI-generated reports using both natural language generation (NLG) and clinical efficacy (CE) metrics. For NLG metrics, which measure linguistic similarities between AI- and radiologist-generated reports, we use BLEU-n (B-n), METEOR (MTR), and ROUGE-L (R-L). For CE metrics, we assess clinical relevance by computing the micro-average Precision (P), Recall (R), and F1-score (F1) across 14 observations annotated by CheXpert (irvin-chexpert).
Implementation details. The unified feature dimension is set to $d=768$, the number of latents to $N=128$, the maximum generation length to $K=100$ tokens, and the beam size to 3 during inference. We apply early stopping and learning rate scheduling for training stability. Additional implementation details are in Appendix Section A.
@@ 4.2 Main Results
Baseline approaches. We compare our PriorRG against 12 state-of-the-art (SOTA) methods, categorized into eight groups: knowledge graph-based methods (KiUT (huang-kiut) and METransformer (wang2023metransformer)), contrastive learning framework (CoFE (cofe-eccv-24)), retrieval-augmented method (DCG (liang2024divide-acmmm-24)), memory alignment approaches (CMN (chen-etal-2021-cross-modal) and MAN (shen2024automatic_aaai)), LLM-based approaches (R2GenGPT (wang-2023-r2gengpt), Med-LLM (liu2024in-context-acmmm), and R2-LLM (aaai-liu2024bootstrapping-llm)), clinical context-driven model (SEI (sei)), longitudinal data-based method (HERGen (2024-eccv-hergen)), and human preference optimization (MPO (aaai-2025-mpo)).
Model Stage 1 Stage 2 NLG Metrics $\uparrow$ CE Metrics $\uparrow$
$C$ PI Hidden States B-1 B-2 B-3 B-4 MTR R-L P R F1
(a) 0.355 0.219 0.149 0.108 0.154 0.262 0.521 0.431 0.472
(b) 0.357 0.216 0.143 0.102 0.154 0.255 0.513 0.480 0.496
(c) 0.405 0.283 0.214 0.170 0.186 0.320 0.517 0.460 0.487
(d) 0.404 0.285 0.217 0.173 0.187 0.328 0.540 0.477 0.507
(e) 0.400 0.282 0.214 0.171 0.186 0.323 0.539 0.465 0.499
(f) 0.387 0.271 0.206 0.165 0.180 0.320 0.526 0.408 0.459
PriorRG 0.412 0.290 0.220 0.175 0.189 0.324 0.541 0.485 0.511
Comparison with SOTA methods. Table 1 presents a comparison between our PriorRG and SOTA methods on MIMIC-CXR (johnson-mimic-cxr-jpg) and MIMIC-ABN (mimic-abn-ori) datasets. The results show that our PriorRG consistently surpasses existing methods in both linguistic quality and clinical accuracy. On the MIMIC-CXR dataset, PriorRG achieves a notable 3.6% improvement in B-4, indicating better matching accuracy in longer n-gram sequences. Additionally, gains in MTR (+1.3%) and R-L (+1.5%) highlight improvements in lexical diversity, synonym matching, and semantic consistency. In terms of CE metrics, PriorRG boosts the F1 by 3.8%, demonstrating greater reliability in clinical accuracy. On the MIMIC-ABN dataset, PriorRG achieves a 5.9% increase in B-1 and a 1.1% improvement in F1, emphasizing its effectiveness in describing abnormal findings. These improvements stem from our proposed prior-aware coarse-to-fine decoding, which enables the model to recognize individual clinical contexts and capture disease progression. Appendix presents extra results (Table 6) and comparisons with MLRG (Tables 79).
Clinical accuracy of 14 observations. Appendix Table 10 presents the clinical accuracy of 14 observations extracted by the CheXpert (irvin-chexpert). Our PriorRG outperforms the clinically-informed SEI model (sei) on 13 of 14 observations in terms of F1-score, highlighting the superior clinical accuracy of generated reports.
Evaluation with large language model. We assess clinical consistency and expert alignment of generated reports using two metrics: “#Matched Findings” and the GREEN score (2024-green). The GREEN score combines clinical significance errors with “#Matched Findings,” offering a robust evaluation aligned with expert preferences. Both metrics are derived via the pre-trained GREEN-RadLlama2-7B model. We benchmark PriorRG against CMN (chen-etal-2021-cross-modal), CGPT2 (nicolson-improving-cvt2distilgpt2), SEI (sei), and MLRG (mlrg). Figure 4 shows that PriorRG significantly outperforms all baselines across both metrics, confirming its ability to generate clinically reliable, expert-aligned reports.
4.3 Ablation Study
Effect of prior-guided contrastive pre-training (Stage 1). We evaluate Stage 1 using a medical image-text retrieval task on the MIMIC-5x200 dataset (johnson-mimic-cxr-jpg). Following the protocols in (pmlr-v182-zhang22a-5x200; nips-2024-benchx), we construct the MIMIC-5x200 dataset by sampling 200 test set instances for each of five common diseases— Atelectasis, Cardiomegaly, Consolidation, Edema, and Pleural Effusion—from MIMIC-CXR, resulting in a total of 1,000 samples. Given a query image, we retrieve the top-K most similar reports from this dataset and evaluate performance using Category-Precision@K (Cat-P@K) and Study-Precision@K (Stu-P@K). Cat-P@K measures whether retrieved reports belong to the same disease category as the query image, while Stu-P@K evaluates whether retrieved reports originate from the same study. As illustrated in Figure 5, PriorRG surpasses BiomedCLIP (zhang-2023-biomedclip), BioViL-T (2023-cvpr-biovit-2301), and ablated variants across both metrics. These results indicate that Stage 1 effectively encodes prior knowledge into semantically rich multimodal representations. Furthermore, Table 2 shows that removing Stage 1 (variant (f)) leads to significant drops in both clinical accuracy and language quality, confirming its central role in improving downstream report generation.
Figure 6: Qualitative comparison of reports from the baseline (sei), our proposed PriorRG, and radiologists. Sentences in the radiologist-generated report are color-coded to match corresponding descriptions in the AI-generated reports. In PriorRG’s output, correct disease progression patterns and failure descriptions are marked in bold and underlined, respectively.
Model NLG Metrics $\uparrow$ CE Metrics $\uparrow$
B-2 B-4 MTR P R F1
LastOnly 0.278 0.168 0.185 0.554 0.466 0.506
Fine2coarse 0.284 0.172 0.187 0.549 0.482 0.514
PriorRG 0.290 0.175 0.189 0.541 0.485 0.511
Metric w/ PI w/o PI w/ CC w/o CC w/ PK w/o PK
B-2 $\uparrow$ 0.289 0.288 0.294 0.139 0.294 0.140
F1 $\uparrow$ 0.516 0.514 0.508 0.492 0.512 0.497
Qualitative analysis. Figure 6 presents a comparative visualization of generated reports on the MIMIC-CXR dataset. In AI-generated reports, more diverse colors indicate broader coverage of clinical findings, while longer bars reflect more accurate and detailed descriptions. The results show that: (1) PriorRG generates high-quality drafts requiring minimal revision—for example, in Case 1, only a brief note on pleural effusion is added by the radiologist. (2) PriorRG captures disease progression effectively, as seen in the correct description of unchanged cardiomegaly in Case 1. (3) PriorRG responds to clinical context—e.g., in Case 2, the model correctly addresses concern about Dobbhoff tube placement and identifies potential complications. These examples demonstrate PriorRG’s ability to track disease progression and understand clinical intent. Additional examples and failure case analysis are provided in Appendix B.
Conclusion
We introduced PriorRG, a chest X-ray report generation framework that leverages clinical priors to enhance both image-text alignment and report quality. Specifically, we proposed a prior-guided contrastive pre-training scheme that mirrors diagnostic reasoning by using clinical context to guide spatiotemporal features extraction. Next, we presented a prior-aware coarse-to-fine decoding that progressively integrates clinical context, disease progression patterns, and hierarchical visual cues, enhancing the clinical accuracy and fluency of generated reports. Extensive experiments validated the effectiveness of our PriorRG on medical image-text retrieval and report generation. Future work will explore the organ-aware diagnosis framework (Gu_2024_WACV-organ-mask) to further enhance interpretability.
Acknowledgments
The work was jointly supported by the National Natural Science Foundations of China [grant number: 62272364], the Provincial Key Research and Development Program of Shaanxi [grant number: 2024GH-ZDXM-47], the Fundamental Research Funds for the Central Universities and the Innovation Fund of Xidian University [grant number: YJSJ25012].
[…]
Prior Image (PI) 0% 0.287 0.513
10% 0.287 0.513
30% 0.288 0.513
50% 0.289 0.515
100% 0.289 0.515
Clinical Context (CC) 0% 0.139 0.492
10% 0.157 0.495
30% 0.187 0.498
50% 0.220 0.501
100% 0.294 0.507
Table 5: Report generation performance on the MIMIC-CXR dataset under varying proportions of Prior Image (PI) and Clinical Context ($C$). The experiment moves beyond the binary setting of Table 4 by gradually adjusting the availability of prior knowledge.
Appendix A Implementation Details
Model architecture. The image encoder (RAD-DINO) (2024-rad-dino) remains frozen throughout training. In contrast, the text encoder (CXR-BERT) (2022-eccv-cxr-bert), the report generator (DistilGPT2) (Sanh2019DistilBERTAD), projection heads, the attention-enhanced layer fusion network (ALF), and the spatiotemporal fusion network (STF) are fine-tuned during both training stages. All experiments are conducted on a single NVIDIA V100 GPU, with a memory usage of approximately 26 GB.
Training schedule. We adopt a two-stage training pipeline. For the MIMIC-CXR dataset (johnson-mimic-cxr-jpg), the model is trained for 30 epochs in Stage 1 (prior-guided contrastive pre-training) with a batch size of 32, followed by another 30 epochs of fine-tuning in Stage 2 (prior-aware coarse-to-fine decoding) with a batch size of 16. In Stage 1, the learning rate is set to 5e-5. In Stage 2, the report generator is optimized with a learning rate of 5e-5, while all other components are trained with a learning rate of 5e-6. For the MIMIC-ABN dataset (mimic-abn-ori), as a derivative of MIMIC-CXR, the model is initialized with weights pre-trained on MIMIC-CXR and fine-tuned for 30 epochs using a learning rate of 5e-6 and a batch size of 16.
Metrics implementation details. All metrics are implemented using the pycocoevalcap (chen2015microsoft-coco-caption) and f1chexbert (Smit2020_chexbert) libraries. For all metrics, higher values indicate better performance.
Medical image-text retrieval. BioViL-T (2023-cvpr-biovit-2301) and BiomedCLIP (zhang-2023-biomedclip) are implemented using their officially released pre-trained weights and inference pipelines. Our PriorRG model is initialized with weights obtained from Stage 1 pre-training. For all three methods, the medical image-report retrieval is conducted by calculating feature similarity between image and report embeddings, followed by retrieving the top-K most similar reports.
Appendix B Additional Experiments
[…]
Zero-shot Report Generation
R2GenGPT Meta-Radio’23 0.112 0.177
Med-LLM MM’24 0.125 0.185
PriorRG (Ours) - 0.178 0.211
Supervised Report Generation
METransformer CVPR’23 0.172 0.192
MAN AAAI’24 0.170 0.213
DCG MM’24 0.186 0.211
Med-LLM MM’24 0.168 0.209
R2-LLM AAAI’24 0.184 0.208
PriorRG (Ours) - 0.196 0.216
Table 6: Comparison of report generation results on the IU X-ray dataset (demner2016preparing-iu-xray) under zero-shot and supervised settings. Best results are in Bold.
Method M/S Stage 1 Stage 2
Trait #Param Trait #Param
MLRG M DP 191M DP+$C$ 296M
PriorRG S DP+$C$ 239M DP+$C$ 335M
Table 7: Comparison between our PriorRG and the recent baseline MLRG (mlrg) in terms of input requirements and model size. “M/S” denotes multi-view or single-view current images; “CC” and “DP” indicate whether a method incorporates clinical context and disease progression, respectively. “#Param” refers to the total number of model parameters.
Method Stu-P@1 Stu-P@3 Stu-P@5 Cat-P@1 Cat-P@3 Cat-P@5
MLRG 0.096 0.096 0.088 0.450 0.448 0.437
PriorRG w/o PK 0.279 0.238 0.191 0.601 0.574 0.549
PriorRG w/o CC 0.280 0.225 0.173 0.602 0.561 0.526
PriorRG w/o PI 0.393 0.324 0.245 0.637 0.597 0.548
PriorRG (Ours) 0.402 0.334 0.251 0.640 0.596 0.544
Method Venu NLG Metrics $\uparrow$ CE Metrics $\uparrow$ Time (s/sample) $\downarrow$ GPU $\downarrow$
B-1 B-2 B-3 B-4 MTR R-L P R F1 RG
Table 9: Performance comparison of report generation on the MIMIC-CXR dataset. Best results are highlighted in Bold. RG denotes F1-RadGraph (jain-radgraph) metric, which evaluates the overlap of clinical entities and relations, and better aligns with radiologists’ assessments (yu2023evaluating).
B.1 Ablation Study on the Contribution of Prior Knowledge
To quantify the contribution of different forms of prior knowledge, we perform an ablation study on the MIMIC-CXR dataset (johnson-mimic-cxr-jpg) by varying the proportions of Prior Image (PI) and Clinical Context (CC) provided to the model. Unlike the binary on/off setup in Table 4, we gradually adjust the availability of each type of prior knowledge to five levels: 0%, 10%, 30%, 50%, and 100%.
As shown in Table 5, both BLEU-2 (B-2) and F1 consistently improve as more prior knowledge becomes accessible. PI shows only marginal influence, and PriorRG maintains stable performance even when PI is largely absent, highlighting its robustness to patient-specific sparsity. In contrast, CC yields a substantially stronger performance gain across all metrics, indicating that contextual clinical cues provide more actionable guidance for report generation. This trend aligns with radiological reasoning, where the integration of patient context is essential for accurate interpretation of imaging findings.
B.2 Comparison with SOTA Methods on the IU X-ray Dataset in Zero-shot and Supervised Settings
Following previous works (chen-etal-2020-generating; chen-etal-2021-cross-modal), we split the IU X-ray dataset (demner2016preparing-iu-xray) into training, validation, and test sets with a 7:1:2 ratio. The detailed distribution is presented in Table 11.
Results in Table 6 show the following: (1) In the zero-shot setting, PriorRG achieves the highest BLEU-4 (0.178) and METEOR (0.211), clearly outperforming baselines like Med-LLM (liu2024in-context-acmmm) and R2GenGPT (wang-2023-r2gengpt). (2) In the supervised setting, PriorRG again achieves the best results, with BLEU-4 of 0.196 and METEOR of 0.216, outperforming strong baselines including METransformer (wang2023metransformer), MAN (shen2024automatic_aaai), DCG (liang2024divide-acmmm-24), Med-LLM (liu2024in-context-acmmm), and R2-LLM (aaai-liu2024bootstrapping-llm). (3) These results demonstrate the strong generalization capability of our PriorRG across datasets and its robustness, even when prior images are unavailable.
B.3 Comparison with Concurrent Work MLRG on the MIMIC-CXR Dataset
MLRG (mlrg), a concurrent work accepted at CVPR 2025, also jointly models temporal visual changes and clinical context. To enable a comprehensive comparison, we evaluate both methods in terms of methodological design (Table 7), medical image-text retrieval (Table 8), and report generation (Table 9). (1) Methodology: Unlike MLRG (mlrg), our PriorRG explicitly models disease progression and clinical context during cross-modal alignment (Stage 1; see Table 7). This design leads to significantly improved retrieval performance, as reflected by higher Stu-P@K and Cat-P@K scores (Table 8). (2) Report Generation: On the chest X-ray report generation task, PriorRG consistently outperforms MLRG across natural language generation (NLG) and clinical efficacy (CE) metrics, achieving a +1.7% improvement in BLEU-4 (B-4) and +0.6% in F1-CheXbert (F1). (3) Efficiency and Deployment: PriorRG requires only a single-view current image while delivering higher-quality reports. It achieves over 2× faster inference and reduces GPU memory usage by nearly 5 GB. All comparisons are conducted on the MIMIC-CXR test set (3,852 samples) under identical hardware, batch size, and num_worker configurations, demonstrating PriorRG’s practical efficiency and deployment readiness.
Figure 7: Qualitative comparison of reports from the baseline (sei), our proposed PriorRG, and radiologists. Sentences in the radiologist-generated report are color-coded to match corresponding descriptions in the AI-generated reports. In PriorRG’s output, correct disease progression patterns and failure descriptions are marked in bold and underlined, respectively.
Observation % SEI (sei) PriorRG (Ours)
@@ B.3 Comparison with Concurrent Work MLRG on the MIMIC-CXR Dataset
Table 11: Statistics of the MIMIC-CXR (johnson-mimic-cxr-jpg), MIMIC-ABN (mimic-abn-ori), and IU X-ray (demner2016preparing-iu-xray) for training, validation, and test sets. “#Image”, “#Sample”, and “#Report” denote the number of medical images, samples, and radiology reports, respectively. “%PI”, “%Indication”, and “%History” represent the proportions of prior images, indications, and medical history, respectively.
B.4 Qualitative Analysis